OpenStudy (anonymous):

If a min force of 500N is req to move a box of mass 100kg resting on horizontal floor the coefficient of friction b/w the box and the floor is (g=10m/s^2)

OpenStudy (anonymous):

I got U = 1/2.........But the answer should be 1/sqrt3

OpenStudy (anonymous):

@Mikael @RaphaelFilgueiras @experimentX

OpenStudy (anonymous):

@sami-21

OpenStudy (anonymous):

500=m*g*U u=500/(100*10)

hartnn (hartnn):

if u write N=U.m.g.sin 60,then u get U=1/(sqrt 3),but i don't know how we could get sin 60 there.....

OpenStudy (anonymous):

i think...may be the answer Given is wrong..))